# Physical models of embryonic epithelial healing: A review

### v1

@@ Discussion
The number of publications on this subject have gradually increased. The number of publications peaked between 2019 and 2020 with publication rates being relatively stable afterwards. This trend suggests an increasing interest within the physics community to address the problem of embryonic healing through modeling approaches, but given the relatively small number of articles this may not be the full picture.
We also looked at the distribution of models covered by the collected articles, where we observe some trends. Prior to 2015, continuum models were more prevalent than cell-based models and data-driven models combined. However, in the past five years, there has been a growing prevalence of cell-based applications towards embryonic healing relative to continuum models.
Another interesting observation is the higher presence of hybrid and data-driven models. These trends are likely a reflection of a broader trend in modeling tissue dynamics. This shift likely reflects an increase in single-cell data availability as well an increased accessibility of high-performance computing resources. While these reasons are framed in the context of general tissue dynamics, they are also applicable to embryonic healing in particular.
@@
Evidence both cell-based and continuum models confirms that efficient embryonic wound closure requires the dynamic coordination between actomyosin purse-string contraction and lamellipodial crawling. Vertex and lattice models show that actomyosin cables dominates in small wounds, while protrusive activity becomes dominant as gaps widen, a switch controlled by adhesiveness, contractility and local curvature [85, 22, 114, 115, 118, 159]. Network models further capture spatial heterogeneity and anisotropic force distributions around irregular wounds, to promote more efficient closure [117], a feature harder to realize in purely continuum models. Notably, several models reproduce the experimentally observed curvature-dependent mode-switch [184, 22, 114, 185, 161]. Lattice models, however struggle to resolve fine-grained geometric features beyond areas and perimeters, limiting their ability to predict curvature-driven feedback on actomyosin dynamics.
@@
Beyond acto-myosin contractility, other cytoskeletal elements, particularly microtubules, remain largely underrepresented in current models, despite their known roles in cell polarity and intracellular transport [25, 27]. Incorporating these cytoskeletal components is not just an additional factor, but it is essential for building a comprehensive mechanochemical model of embryonic wound repair.
@@
There is no shortage of modeling efforts addressing key features of embryonic wound repair, yet important gaps persist within certain frameworks. Lattice models, for example, offer fine‑grained control of subcellular mechanics but struggle to represent evolving wound geometry, whereas phase‑field schemes capture smooth cell-boundary deformations yet omit explicit junctional and cytoskeletal dynamics. Hydrodynamic approaches readily incorporate polarity fields, chemical gradients, and large‑scale flow, making them ideal for wounds much larger than a single cell, but their continuum assumptions wash out the heterogeneity and anisotropy intrinsic to real tissues. Recognizing these trade‑offs reveals opportunities for extending each framework: lattice models could adopt curvature‑sensitive rules to better handle shape, phase fields might integrate discrete junctional forces, and hydrodynamic theories can borrow heterogeneity from network approaches.
@@
Looking ahead, we expect hybrid models to become increasingly central, particularly in integrating ECM-tissue mechanics, incorporating underexplored cytosketelal elements and chemical signaling pathways. As hybrid models become more complex, the need for data-driven parameter inference will become more acute. Their greatest limitation, the lack of robust coupling parameters, could be overcome through tighter integration with experimental data and machine learning pipelines.
While individual models offer focused results on specific aspects, the convergence of findings across different methods underscores a broader principle: mechanical forces and chemical signaling are intrinsically linked in embryonic healing. Models that successfully integrate both, particularly through hybrid or data-driven approaches, are best positioned to advance the field.

### v2

@@ Discussion
The models considered in this review can be classified by how tissue behavior emerges. Cell-based models are rule-based, with defined behaviors for the individual cells that lead to emergent properties at the tissue level. Continuum models can be phenomenological to varying degrees, depending on the framework. Viscoelastic models and the standard phase field models are phenomenological, as the properties expected are directly incorporated into the energy functional. In contrast, hydrodynamic models are derived from physical principles, and it is expected that tissues match their behavior in some conditions.
@@
Evidence across both cell-based and continuum models confirms that efficient embryonic wound closure requires the dynamic coordination between actomyosin purse-string contraction and lamellipodial crawling. Vertex and lattice models show that actomyosin cables dominates in small wounds, while protrusive activity becomes dominant as gaps widen, a switch controlled by adhesiveness, contractility and local curvature [88, 22, 117, 118, 121, 162]. Network models further capture spatial heterogeneity and anisotropic force distributions around irregular wounds, to promote more efficient closure [120], a feature harder to realize in purely continuum models. Notably, several models reproduce the experimentally observed curvature-dependent mode-switch [187, 22, 117, 188, 164]. Lattice models, however struggle to resolve fine-grained geometric features beyond areas and perimeters, limiting their ability to predict curvature-driven feedback on actomyosin dynamics.
@@
Beyond acto-myosin contractility, other cytoskeletal elements, particularly microtubules and intermediate filaments, remain largely underrepresented in current models, despite their known roles in cell polarity and intracellular transport [25, 27]. Additionally, the influence of the nucleus must be considered as it is not often considered in wound healing models. Incorporating these cytoskeletal components is not just an additional factor, but it is essential for building a comprehensive mechanochemical model of embryonic wound repair.
@@
There is no shortage of modeling efforts addressing key features of embryonic wound repair, yet important gaps persist within certain frameworks. Lattice models, for example, offer fine‑grained control of subcellular mechanics but struggle to represent evolving wound geometry and lack true dynamics, whereas phase‑field schemes capture smooth cell-boundary deformations yet omit explicit junctional and cytoskeletal dynamics. Hydrodynamic approaches readily incorporate polarity fields, chemical gradients, and large‑scale flow, making them ideal for wounds much larger than a single cell, but their continuum assumptions wash out the heterogeneity and anisotropy intrinsic to real tissues. Recognizing these trade‑offs reveals opportunities for extending each framework: lattice models could adopt curvature‑sensitive rules to better handle shape, phase fields might integrate discrete junctional forces, and hydrodynamic theories can borrow heterogeneity from network approaches.
The choice of the model also depends on the intended purpose. Simpler models with fewer equations or parameters allow general principles to be inferred and are more interpretable, but may be less quantitatively accurate. More detailed models can reproduce observed tissue behavior more faithfully, though at the cost of interpretability. The appropriate framework depends on whether the goal is to understand underlying mechanisms or to closely match experimental observations.
Depending on the biological scale and processes of interest, different models are appropriate. Where individual cell morphologies are relevant, lattice or multiphase field models are adequate. Vertex models are well suited for junctional mechanics, while hydrodynamics or viscoelastic models capture tissue-scale collective behavior. Wound geometry also affects which mechanisms dominate: purse-string forces are negligible in scratch assays, whereas crawling is less prominent in non-adherent substrates. This can guide model selection for experimentalists studying epithelial wound healing.
@@
Other forms of tissue repair, like single-cell extrusions, which occur in three dimensions, or transient junctional tears, constitute the most common wounds in embryonic epithelia. These wounds are treated as topological defects rather than free boundaries in a plane, as in the models considered here. While biologically relevant, these processes fall outside the scope of this review, which focuses on wounds dominated by collective, multicellular dynamics. There are, however, some recent examples of physical models applied to study cell-extrusions and junctional tears [108, 257, 258, 259]. Although these processes are physically distinct from multicellular free-boundary wounds, they share underlying mechanisms, which shows the broad range of epithelial repair mechanisms.
While this review focuses on embryonic epithelial healing, many of the physical results obtained from these models are relevant to adult wound repair, provided the underlying assumptions are reconsidered carefully. Model implementations related to collective cell migration, proliferation and mechanochemical feedback are largely transferable, although their results must be re-evaluated in three dimensions. In particular, reaction-diffusion models for signaling, or active matter models to describe polarity can be extended to adult healing with some formal changes.
The major differences are in tissue architecture and mechanical environment. While embryonic healing is three dimensional, it can often be well described by two dimensional models as we have seen across this review. Adult wound healing however occurs in a three dimensional extracellular matrix populated by fibroblasts and being actively remodelled. In contrast, in the embryonic models reviewed here, ECM is a quasi-two-dimensional substrate supporting a confluent epithelial sheet. ECM is a material that experiences plastic deformations and fiber realignment, and plays a far more active role in adult wound healing. Thus the models which include ECM as an elastic or passive surface in embryonic healing contexts are not directly applicable.
Additional processes central in adult healing, such as inflammation and angiogenesis, introduce new active components and length scales. While these processes lie outside the scope of the present review, it is notable that several of the models discussed here, have already been adapted to describe angiogenic growth and vascular remodelling, where three dimensional geometry is essential and cannot be abstracted away [254, 260].
@@
Looking ahead, we expect hybrid models to become increasingly central, particularly in integrating ECM-tissue mechanics, incorporating underexplored cytosketelal elements and chemical signaling pathways. As hybrid models become more complex, the need for data-driven parameter inference will become more acute. Their greatest limitation, the lack of robust coupling parameters, could be overcome through tighter integration with experimental data and machine learning pipelines. To contextualize these modeling directions, Table 3 summarizes the experimental systems, species, wounding assays, and open-source implementations associated with the models reviewed. The overview is not exhaustive; the absence of specific assays or biological systems reflects the current literature and not intrinsic limitation of the models. It represents an opportunity for future exploration.
Model Biological System Species Wounding Assay Open-source implementations
Lattice Epithelial monolayer (in vitro) Human (OVCAR-3, MCF-7); Canine (MDCK) Scratch assay; Laser ablation; Barrier removal assay CompuCell3D; Morpheus
Network Embryonic epithelial monolayer (in vivo); embryonic-like monolayer (in vitro) Canine (MDCK); Drosophila embryo Laser ablation; Barrier removal assay SurfaceEvolver; SAMoS; cellGPU
Phase-field Embryonic-like epithelial monolayer (in vitro) Canine (MDCK) Scratch assay; Barrier removal assay Finite element solvers (FEniCS, FreeFEM)
Hydrodynamic Epithelial monolayer (in vitro) Canine (MDCK); Human keratinocytes Scratch assay; Barrier removal assay Finite element solvers (FEniCS, FreeFEM)
Viscoelastic Embryonic epithelial tissue (in vivo); epithelial monolayer (in vitro) Canine (MDCK); Human keratinocytes; Chick embryo Incision; Barrier removal assay Finite element solvers (FEniCS, FreeFEM)
Table 3: Summary of the biological systems, species, wounding assays, and open-source implementations for the main physical models of epithelial wound healing reviewed. The table distinguishes between in vivo embryonic tissue and in vitro monolayer systems, highlighting where the reviewed models have been applied. Examples listed are not exhaustive. Hybrid and Data-driven systems are not included, as their application is not restricted to specific systems.
While individual models offer focused results on specific aspects, the convergence of findings across different methods underscores a broader principle: mechanical forces and chemical signaling are intrinsically linked in embryonic healing. At present, however, this convergence is primarily assessed qualitatively, based on whether models reproduce similar experimentally observed behaviors. Models that successfully integrate both, particularly through hybrid or data-driven approaches, are best positioned to advance the field. A more systematic comparison would require deriving effective continuum parameters from cell-based models, enabling quantitative links between microscopic rules and macroscopic material properties. Such a formal theoretical translation between models remains to be developed and constitutes an important open challenge.